# CREST-Search: Comprehensive Red-teaming for Evaluating Safety Threats in Large Language Models Powered by Web Search

### v1

@@ 3.1. Problem Statement
(1) $R(q)\rightarrow D=\{d_{1},d_{2},...,d_{k}\}$
where $D$ is a set of $k$ webpages or document snippets returned by the search engine. The LLM then conditions on both the original query and retrieved evidence to generate the final response:
(2) $M(q,D)\rightarrow r$
where $r$ is the LLM’s answer, containing references from $D$.
According to Luo et al. (Luo et al., 2025), safety risks associated with search-enabled LLMs can be categorized into the following three levels:
@@
    Response risk. The response generated by the target LLM contains harmful content, such as harassment or other forms of toxicity.
@@
    Citation risk. The LLM cites external web links that contain harmful content, including the modalities of text or images.
@@
    Combined risk. Both the LLM response and cited webpages include harmful content, compounding the overall severity.
@@
A red-teaming methodology should satisfy three core requirements. (1) It must reliably expose real safety failures across diverse target models and deployment settings. (2) It should produce a broad and representative set of test cases that cover as many harmful content categories and attack modalities as possible. (3) The generation process must be efficient and cost-effective, imposing minimal computational, time, and financial overhead to enable practical, large-scale evaluation.
3.2. Threat Model
We adopt a standard threat model in which a tester seeks to assess the security of an LLM system augmented with web search. Such red-teaming may be performed under several practical scenarios: (1) a model developer conducts internal testing to identify and remediate vulnerabilities prior to public release; (2) a prospective customer evaluates competing LLM services to select the one that best meets their safety and functional requirements; or (3) a trusted third party is engaged to perform audits and issue safety certifications. In all cases, the tester’s objective is to simulate realistic adversarial conditions against the complete search-based LLM pipeline, revealing end-to-end failures that could arise in real-world operation.
@@ 3.2. Threat Model
To ensure broad applicability, we adopt a black-box setting, where the tester has no internal knowledge of the target LLM system, including (1) the model architecture and parameters; (2) the retrieval and search-integration mechanisms; (3) any deployed safety filters or mitigation strategies. The tester’s interaction is therefore limited to externally observable behavior: he may only craft and submit prompts to the system and inspect the returned responses. Under this constraint, all evaluations must be performed using input–output probing alone, which reflects realistic auditing scenarios faced by developers, customers, and third-party auditors. We further discuss the ethical considerations of conducting such adversarial testing in Appendix A.1, ensuring that our evaluation remains responsible and compliant.
4. Methodology
To comprehensively evaluate the safety of LLMs with web search, we present our red-teaming approach, CREST-Search, as shown in Figure 1. It adopts a three-stage pipeline to automatically generate effective test queries against the target system (Section 4.1). We further introduce a novel adversarial search query generation strategy (Section 4.2) and red-teaming model construction approach (Section 4.3), to enhance the quality of the generated queries.
4.1. Red-teaming Pipeline
CREST-Search introduces an execution pipeline, to iteratively refine and produce effective testing queries against a given search-enabled LLM. This pipeline comprises three main stages: adversarial search query generation, web search execution and risk evaluation, and adversarial search query refinement
1. Adversarial search query generation. A red-teaming model is introduced to guide the automatic generation of adversarial search queries following distinct strategies. These queries can trigger the target LLM to output various categories of malicious content, exposing the citation risk mentioned in Section 3.1. The details of the generation strategies are described in Section 4.2. To be specific, we design a structured framework that combines a specialized system prompt and a user prompt with explicit guidance. The system prompt requires the red-teaming model to generate queries following the fundamental rule that the adversarial search queries appear benign but retrieve and cite webpages containing harmful content by effectively exploiting the search mechanism. The user prompt provides the LLM with concrete instructions to generate the query that retrieves a specific harmful content category under a certain generation strategy. Another requirement is that the final query set should retain diversity, keeping each query distinct from others. The detail of the system prompt is provided in Appendix A.4.
2. Web search execution and risk evaluation. Once the adversarial queries are generated, we feed them into the target LLM and evaluate the outcomes by analyzing the external webpages’ content that it cites. This step explicitly reflects the distinct attack surface of search-enabled LLMs, which differs from conventional red-teaming focused only on text generation. We employ two advanced moderation tools, Llama Guard (Inan et al., 2023) and OpenAI Moderation (OpenAI, 2024b), to systematically assess the toxicity of multimodal content in the webpages. An adversarial search query is deemed successful if harmful content is detected in the cited sources. Otherwise, the query enters the refinement loop for optimization.
3. Adversarial search query refinement. If the initial query fails to induce the target LLM to cite harmful webpages, we initiate a refinement process tailored for the web search setting. We employ another LLM as a judgment model for this step by adopting in-context learning (Dong et al., 2022). The judgment model incorporates the original query together with the applied generation strategy and the target harmful content category, enabling it to fully capture the context of the attack scenario. Then it provides feedback to the red-teaming model for revision, which maintains the original attack scenario with the strengthened refinement objective: modifying the query so that the target model retrieves and cites webpages containing harmful content in its response. The refinement process terminates once the query successfully triggers the citation of harmful content or when the maximum number of refinement rounds is reached.
This refinement loop offers two advantages: (1) It makes our framework more effective and efficient in online exposing citation risks, which are often overlooked by conventional red-teaming approaches. (2) It can be applied offline to construct domain-specific adversarial datasets. These datasets enable the fine-tuning of general LLMs into specialized red-teaming models, thereby improving the quality of the initially generated queries. As a result, the need for massive repeated optimization decreases, leading to lower computational costs and shorter optimization time. This makes the framework more practical for large-scale red-teaming deployment. The details of this fine-tuning process are provided in Section 4.3.
4.2. Search Query Generation Strategy
While prior works (Luo et al., 2025; Yi et al., 2024) examine adversarial prompts for vanilla LLMs and their resistance to unsafe generation, they overlook the distinct threat surface of LLMs with web search. The retrieval–citation process can expose users to unsafe or low-credibility content via seemingly benign queries.
@@ 4.2. Search Query Generation Strategy
To address this gap, we propose a novel set of adversarial search query generation strategies tailored to the LLM-with-search scenario. Unlike traditional red-teaming prompts, which primarily test a model’s intrinsic safety filters, our strategies are explicitly designed to exploit vulnerabilities in the retrieval–citation pipeline. Specifically, we adopt three strategies for generating adversarial queries: keyword injection, exaggeration, and role play. Their goal is to shift the risk from direct unsafe generation to unsafe citation, which is more subtle and harder to detect with conventional safeguards. The details of these strategies are described as follows.
@@
    Keyword injection. This strategy selects harmful or sensitive keywords that often exist in malicious websites, and deliberately inserts them into otherwise benign queries, aiming to steer the retrieval module of the system toward unsafe webpages. By embedding such keywords that resemble the search-engine optimization (SEO) bait (Yalçın and Köse, 2010), the target LLM is more likely to retrieve and cite malicious or low-credibility sites, thereby exposing users to harmful web content.
@@
    Role play. This strategy guides the LLM to adopt a specific identity, context, or scenario that conceals the malicious intent of the query. For example, instructing the model to “act as a legal advisor” or “simulate a cybersecurity analyst” may bypass its safety filters/mechanisms and cause it to fetch and cite harmful sites as seemingly authoritative evidence.
4.3. Red-teaming Model Construction
@@ 4.3.1. WebSearch-Harm Dataset
To prepare high-quality training data for red-teaming model fine-tuning, we construct an adversarial query dataset called WebSearch-Harm, covering all the combinations of harmful content categories and generation strategies described in Section 4.1. The construction process comprises four stages: seed adversarial search query generation, adversarial search query refinement, query set preprocessing, and dataset evaluation. To enhance the initial quality of seed adversarial queries, beyond fully exploiting the iteration loop, we introduce a novel two-stage few-shot expansion strategy. Inspired by few-shot learning (Wang et al., 2020), this strategy provides a principled way to optimize seed initialization, replacing the random approach. In the first stage, we create a small seed set of few-shot exemplars manually written by human experts. These handcrafted queries are designed based on experts’ knowledge of the LLM-with-web-search scenario, integrating insights from both conventional LLMs and search engine mechanisms. After executing and refining this initial seed set, we retain only those queries that successfully trigger risks as optimized cases. In the second stage, these optimized adversarial queries are reused as few-shot exemplars to guide the model in generating a larger pool of seed adversarial queries continuously. This stage replaces purely empirical expert-written examples with practically validated cases, enabling the model to better learn adversarial patterns and improve efficiency. A subsequent round of execution and refinement is then performed to further consolidate quality. This two-stage expansion strategy allows the LLM to benefit from expert knowledge and validated adversarial exemplars, thereby accelerating convergence and improving the effectiveness of the refinement process.
@@
Then, we conduct data preprocessing to improve the dataset quality by filtering out duplicates and removing the failure cases that cannot trigger the target LLM to retrieve the harmful content. Each training instance is annotated with an input instruction specifying the harmful content category and the query construction strategy, paired with the corresponding adversarial query that successfully triggers risky outputs in the target LLM. Finally, we carry out a dataset evaluation, including statistical analyses of dataset size, the coverage of harmful categories, and the distribution of construction strategies. When imbalances are detected, we selectively augment underrepresented types to achieve a more uniform distribution. The WebSearch-Harm dataset provides a robust foundation for constructing the red-teaming model. The detailed distribution and analysis are described in Appendix A.2.
4.3.2. Supervised Fine-tuning for Web Search
@@
We adopt supervised fine-tuning (SFT) (Dong et al., 2023) to train a generic LLM into a specialized adversarial query generator. Benefit from the curated adversarial search pairs in WebSearch-Harm, the model acquires the ability to generate more realistic, diverse, and potent adversarial search queries, thereby improving the overall scalability and consistency of our red-teaming framework.
5. Experiments
5.1. Setup
@@
Target models. We select four representative LLMs with web search functions to evaluate the effectiveness of our proposed red-teaming approach: GPT-4o-search-preview, GPT-4o-mini-search-preview, Gemini-2.0-flash-search, Gemini-2.5-flash-search.
Baselines. We compare with 4 relevant baselines. (1) UAT (Zou et al., 2023): Universal and transferable adversarial attacks on aligned language models; (2) HarmfulQA (Bhardwaj and Poria, 2023): Chain of Utterances-based (CoU) prompting jailbreak includes harmful questions; (3) DangerousQA (Bhardwaj and Poria, 2023): Chain of Utterances-based (CoU) prompting jailbreak includes dangerous questions; (4) HarmBench (Mazeika et al., 2024): A publicly available dataset for evaluating harmful outputs in standard offline LLMs.
@@
Method configuration. CREST-Search employs multiple LLMs for different roles. During the dataset construction phase for the model fine-tuning, we employ the refinement loop. In this phase, we employ GPT-4o-mini as the test case generation model, GPT-4o-search-preview as the target victim model and GPT-4o as the refinement model. Each generation–optimization loop is executed for at most 10 rounds. Once the harmful content is detected in any cited webpages, we skip to the next case for optimization. Then we set Gemini-2.5-flash as the base fine-tuning model.
@@
    Risk detection rate. (RDR) It evaluates the proportion of queries that successfully trigger the risks mentioned in Section 3.1. We use 2 detectors (OpenAI, 2024b; Fedorov et al., 2024) to evaluate the harmful content in the response or websites. If any one of the detectors predicts that the response/citation is risky, we consider the query to have successfully triggered the corresponding risk of the LLMs.
@@
Experiment setup. To evaluate the performance of CREST-Search, we employ it to generate a corpus of test cases and execute these adversarial queries on the target victim models directly without any system prompt settings. To achieve stable experiment results, we run each query 3 times. Concretely, we consider 5 harmful-content categories and 3 generation strategies mentioned in Section 4.1, and we produce 100 test cases per (category, strategy) pair, total of 1,500 test cases per experimental run.
5.2. Results
In this section, we first compare the performance of CREST-Search with other baselines and detail all types of risks. Next, we analyze the effectiveness of the query generation strategy and the transferability of CREST-Search. Finally, we conduct an ablation study to validate the contributions of each component.
Risk analysis. The red-teaming performance comparison between CREST-Search and baselines is given in Table 1. CREST-Search substantially outperforms all baselines, achieving a risk detection rate of 80.5%, whereas the best-performing baseline (HarmBench) only reaches 11.6%. Moreover, CREST-Search maintains the lowest query toxicity (23.6%), far below that of test cases generated by other baselines. In terms of query diversity, CREST-Search also leads with a score of 0.82. Figure 2 provides a detailed distribution of the total risks triggered by baseline methods and CREST-Search, including response risk, citation risk, and combined risk. The results reveal that the vast majority of risks uncovered by CREST-Search are citation risks, with a proportion of 89.3% among the total risk. Importantly, we observe that many cases involve only citation risks while the generated responses themselves remain benign, with a proportion of 74.7% among the total risk, making them particularly imperceptible and dangerous. Such vulnerabilities are invisible to conventional red-teaming approaches that focus solely on unsafe generation, but are critical in LLMs with web search, where harmful content may originate from external citations.

### v2

@@ 3.1 Problem Statement
$R(q)\rightarrow D=\{d_{1},d_{2},...,d_{k}\}$ (1)
where $D$ is a set of $k$ webpages or document snippets returned by the search engine. The LLM then synthesizes answers using the original query and search results:
$M(q,D)\rightarrow r$ (2)
where $r$ is LLM’s answer, containing references from $D$. According to Luo et al. (2025), safety risks associate with search-enabled LLMs can be categorized into the following three levels:
@@
    Response risk. The LLM response contains harmful content, such as harassment or other forms of toxicity.
@@
    Citation risk. Cited external web links contain harmful content, including modalities of text or images.
@@
    Combined risk. Both LLM response and cited webpages include harmful content, compounding overall severity.
@@
A red-teaming solution should satisfy three core requirements. (1) It must reliably expose real safety failures across diverse target models and deployment settings. (2) It should produce a broad and representative set of test cases that cover as many harmful content categories and attack modalities as possible. (3) The generation process must be cost-effective, imposing minimal computation, time, and financial overhead to enable practical, large-scale evaluation.
3.2 Threat Model
@@
We consider a standard red-teaming threat model in which a tester evaluates the safety of an LLM system augmented with web search by simulating realistic adversarial conditions. The goal is to uncover end-to-end safety failures that may arise during real-world deployment of search-enabled LLM services. We adopt a black-box setting, where the tester has no access to internal system details (e.g., model architecture, retrieval mechanisms, deployed safety filters) and external webpages. The tester can only craft and submit prompts to the system and inspect the returned responses. Under this constraint, all evaluations must be performed using input–output probing alone, reflecting realistic auditing scenarios faced by developers or third-party auditors.
4 Methodology
We present CREST-Search, a red-teaming framework to comprehensively evaluate the safety of LLMs equipped with web search. Its overview is shown in Figure 2. We first introduce a set of adversarial search query generation strategies, which define how to construct queries to induce unsafe retrieval and citation behaviors (§4.1). Building upon these strategies, we design an automated pipeline for adversarial search queries generation, evaluation, and refinement under realistic black-box constraints (§4.2). Finally, we describe a red-teaming model construction approach to strengthen query generation quality (§4.3).
4.1 Search Query Generation Strategy
@@
We propose a novel set of adversarial search query generation strategies tailored to the LLM-with-search scenario. Unlike traditional red-teaming prompts, which primarily test a model’s intrinsic safety filters, our strategies are explicitly designed to exploit vulnerabilities in the retrieval–citation pipeline. Their goal is to shift the risk from direct unsafe generation to unsafe citation, which is more subtle and harder to detect with conventional safeguards. Specifically, we adopt three strategies for generating adversarial queries.
@@
    Keyword injection. This strategy selects harmful or sensitive keywords that often exist in malicious websites, and deliberately inserts them into otherwise benign queries, aiming to steer the retrieval module of the system toward unsafe webpages. By embedding such keywords that resemble the search-engine optimization (SEO) bait (Yalçın and Köse, 2010), the target LLM is more likely to retrieve and cite malicious or low-credibility sites.
@@
    Role play. This strategy guides the LLM to adopt a specific identity, context, or scenario that conceals the malicious search intent of the query. For example, instructing the model to “act as a legal advisor” or “simulate a cybersecurity analyst” may bypass its safety filters/mechanisms and cause it to fetch and cite harmful sites as seemingly authoritative evidence.
Compared to prior red-teaming techniques, these strategies operate under fundamentally different constraint formulations, targeting distinct attack objectives.
4.2 Red-teaming Pipeline
We build an end-to-end red-teaming pipeline, consisting of the following steps.
1. Adversarial Search Query Generation. Guided by the generation instruction, a red-teaming model produces adversarial queries that appear benign yet are likely to induce unsafe citations, covering a diverse set of attack scenarios. Each attack scenario specifies a concrete adversarial objective, which is defined by the combination of an attack strategy and a targeted harmful content category. By combining different attack strategies and harmful categories, CREST-Search systematically explores different mechanisms for exploiting the retrieval–citation pipeline. For each scenario, generated queries are required to (i) remain non-obviously malicious, (ii) retrieve webpages containing harmful or low-credibility content aligned with the target category, and (iii) encourage the target model to cite such content in its response, thereby exposing risks defined in §3.1. We further promote diversity within each attack scenario, encouraging the red-teaming model to generate multiple distinct queries that target the same adversarial objective through varied lexical and semantic realizations, rather than minor revisions. Implementation details of prompt templates are provided in Appendix A.4.
2. Web Search Execution and Risk Evaluation. Given a set of adversarial search queries, we execute them on the target LLM equipped with web search and evaluate the resulting safety outcomes by inspecting the content of the externally cited webpages. Different from conventional red-teaming approaches focused solely on response generation, this step explicitly targets the retrieval–citation attack surface that is unique to search-enabled LLMs. To assess potential safety violations, we apply automated content moderation mechanisms to the cited webpages to identify harmful or toxic content across different modalities. The evaluation results are then forwarded to a judgment model in the subsequent stage to guide the refinement.
3. Adversarial Search Query Refinement. Based on the evaluation results from the previous stage, we perform a judgment-guided refinement process to iteratively improve adversarial search queries. A judgment model is employed to interpret the evaluation results, considering three types of risk outcomes: response risk, citation risk, and their combination. Citation risk serves as the primary criterion for determining attack success. Meanwhile, response risk is incorporated as an auxiliary diagnostic signal to help the judgment model distinguish different failure modes.
If the query is deemed unsuccessful, the judgment model leverages in-context learning techniques (Dong et al., 2022) for refinement guidance generation, incorporating the original adversarial query, the associated attack strategy, the targeted harmful content category, and the observed risks. The guidance suggests how the query should be revised to better induce unsafe citations while preserving the original attack scenario. This guidance is then used by the red-teaming model to update the query. The refinement process iterates until citation risk is triggered or a predefined maximum number of refinement rounds is reached. Upon termination, we further analyze the outcomes and produce comprehensive risk reports.
4.3 Red-teaming Model Construction
@@
WebSearch-Harm Dataset. To prepare high-quality training data for red-teaming model fine-tuning, we construct an adversarial query dataset called WebSearch-Harm, covering all the combinations of harmful content categories and generation strategies described in §4.2. The construction process comprises four stages: seed adversarial search query generation, adversarial search query refinement, query set preprocessing, and dataset evaluation. To enhance the initial quality of seed adversarial queries, beyond fully exploiting the iteration loop, we introduce a novel two-stage few-shot expansion strategy. Inspired by few-shot learning (Wang et al., 2020), this strategy provides a principled way to optimize seed initialization, replacing the random approach. In the first stage, we create a small seed set of few-shot exemplars manually written by human experts. These handcrafted queries are designed based on experts’ knowledge of the LLM-with-web-search scenario, integrating insights from both conventional LLMs and search engine mechanisms. After executing and refining this initial seed set, we retain only those queries that successfully trigger risks as optimized cases. In the second stage, these optimized adversarial queries are reused as few-shot exemplars to guide the model in generating a larger pool of seed adversarial queries continuously. This stage replaces purely empirical expert-written examples with practically validated cases, enabling the model to better learn adversarial patterns and improve efficiency. A subsequent round of execution and refinement is then performed to further consolidate quality. This two-stage expansion strategy allows the LLM to benefit from expert knowledge and validated adversarial exemplars, thereby accelerating convergence and improving the effectiveness of the refinement process.
@@
Then, we conduct data preprocessing to improve the dataset quality by filtering out duplicates and removing the failure cases that cannot trigger the target LLM to retrieve the harmful content. Each training instance is annotated with an input instruction specifying the harmful content category and the query construction strategy, paired with the corresponding adversarial query that successfully triggers risky outputs in the target LLM. Finally, we carry out a dataset evaluation, including statistical analyses of dataset size, the coverage of harmful categories, and the distribution of construction strategies. When imbalances are detected, we selectively augment underrepresented types to achieve a more uniform distribution. The WebSearch-Harm dataset provides a robust foundation for constructing the red-teaming model. The detailed statistics are described in Appendix A.2.
@@
Supervised Fine-tuning for Web Search We adopt supervised fine-tuning (SFT) (Dong et al., 2023) to train a generic LLM into a specialized adversarial query generator. Benefiting from the curated adversarial search pairs in WebSearch-Harm, the model acquires the ability to generate more realistic, diverse, and potent adversarial search queries, thereby improving the overall scalability and consistency of our red-teaming framework.
5 Experiments
5.1 Setup
@@
Target models. We select four representative LLMs with web search to evaluate the effectiveness of our red-teaming approach: GPT-4o-search-preview, GPT-4o-mini-search-preview, Gemini-2.0-flash-search, Gemini-2.5-flash-search.
Baselines. We compare CREST-Search with 4 relevant baselines. (1) UAT (Zou et al., 2023): a gradient-guided iterative discrete optimization method to induce universal and transferable jailbreak behaviors; (2) RED-EVAL (Bhardwaj and Poria, 2023): a conversational red-teaming baseline that models adversarial probing as a multi-turn interaction process to elicit harmful behavior; (3) DangerousQA (Shaikh et al., 2023): Chain of Utterances-based (CoU) prompting jailbreak including dangerous questions; (4) HarmBench (Mazeika et al., 2024): A framework that formulates automated red teaming as an iterative test-case generation process to elicit predefined harmful behaviors from LLMs.
@@
Configuration. CREST-Search employs multiple LLMs for different roles. During the dataset construction phase, we employ GPT-4o-mini as the test case generation model, GPT-4o-search-preview as the target victim model and GPT-4o as the refinement model. Each generation–optimization loop is executed for at most 10 rounds. Then we set Gemini-2.5-flash as the base fine-tuning model. Further deployment and compliance considerations are discussed in Appendix A.1.
We employ CREST-Search to generate a corpus of test cases and execute them on the target victim models directly without any system prompt settings. To achieve stable experiment results, we run each query 3 times. Concretely, we consider 5 harmful-content categories and 3 generation strategies mentioned in §4.2, and produce 100 test cases per (category, strategy) pair, in total of 1,500 test cases per experimental run.
@@
    Risk detection rate. It evaluates the proportion of queries that successfully trigger the risks mentioned in §3.1. We use 2 detectors (OpenAI, 2024b; Fedorov et al., 2024) to evaluate the harmful content in the response or websites. If any one of the detectors predicts the response/citation risky, we consider corresponding risks detected.
@@
5.2 Main Results
Risk analysis. The red-teaming performance comparison between CREST-Search and baselines is given in Table 1. CREST-Search substantially outperforms all baselines, achieving a risk detection rate of 80.5%, whereas the best-performing baseline (HarmBench) only reaches 11.6%. Moreover, CREST-Search maintains the lowest query toxicity (23.6%), far below that of test cases generated by other baselines. In terms of query diversity, CREST-Search ranks second among all methods, achieving a score of 0.59, indicating strong diversity. Although HarmBench achieves the lowest score, it exhibits substantially lower risk detection performance. Figure 3 provides a detailed distribution of the total risks triggered by baseline methods and CREST-Search, including response risks, citation risks, and combined risks. The results reveal that the vast majority of risks uncovered by CREST-Search are citation risks, with a proportion of 89.3% among the total risks. Importantly, we observe that many cases involve only citation risks while the generated responses themselves remain benign, with a proportion of 74.7% among the total risks, making them particularly imperceptible and dangerous. Such vulnerabilities are invisible to conventional red-teaming approaches that focus solely on unsafe generation, but are critical in LLMs with web search, where harmful content may originate from external citations.
Table 1: Performance comparison of baseline models.
   Detection Rate    Toxicity    Diversity
   UAT    5.2%    87.9%    0.77
   RED-EVAL    5.7%    49.9%    0.76
   DangerousQA    10.5%    65.5%    0.61
   HarmBench    11.6%    53.8%    0.48
   CREST-Search    80.5%    23.6%    0.59
Figure 3: Detailed risks analysis across baseline models.